Twenty percent of the 400 students at a certain high school has a family income which is below the poverty line. If 2/5 of these students are females, how many are males?(answer is 48 show your solution)
Answer by Theo(13342)

20% of the students have a family income below the poverty line.
that makes 400 * .2 = 80 students.
2/5 of those are females which means that 3/5 of those are males.
3/5 * 80 = 3*16 = 48